OpenStudy (anonymous):

Given ln(x−5)+ln(x−6)=1, Solve for x

Nnesha (nnesha):

quotient rule\[\large\rm ln y - ln x = \ln \frac{ x }{ y}\] subtraction to division product rule \[\large \rm ln x + \ln y = \ln ( x \times y )\] addition ----> multiplication power rule \[\large\rm ln x^y = y \ln x\]

Nnesha (nnesha):

which one you should apply ?

OpenStudy (anonymous):

Well, what I have so far is: ln(x-5)(x-6) = ln e (x-5)(x-6) = e^1 x^2-11x+30 = e^1 then i'm stuck

Nnesha (nnesha):

http://prntscr.com/76vd3o how did you get ln at right side ?

OpenStudy (anonymous):

I just replaced 1 with ln e thinking that it might be easier that way since ln e = 1 but yeah still stuck

Nnesha (nnesha):